2. A constant force changes the speed of a 100kg sprinter from 4.0m/s to 6.0 m/s in 1.0 s. Calculate
a) The acceleration of the sprinter during this time.
b) The magnitude of the force
1
Expert's answer
2019-09-16T10:02:52-0400

(a) The acceleration of the sprinter


"a=\frac{\Delta v}{\Delta t}=\frac{6.0\:\rm m/s-4.0\: m/s}{1.0\:\rm s}=2.0\:\rm m/s^2"

(b) The magnitude of the force


"F=ma=100\:\rm kg\times 2.0\: m/s^2=200\: N"